### v2

@@ 3.3 Equilibrium Policy Platforms
We can now determine equilibrium platform choices as functions of the parameters of the game. If severe climate change leads to catastrophic baseline welfare level losses, $\Delta>1$, then voters ignore any signal indicating that $\omega=1$. Our analysis so far reveals that then a majority always votes for policy 0, whenever $p_{1}\neq p_{2}$. Proposition 2 tells us that in this case candidates never campaign on policy $p=1$, independent of the information they hold. Hence, if severe climate change is catastrophic, then candidates will ignore any information and always choose the optimal policy for mild climate change, $\mathbf{p}^{*}=(0,0)$.
When $\Delta\leq 1$, the set of equilibria becomes larger. The reason is that now voters’ beliefs about the enacted policy become self-fulfilling. If voters trust the government in the sense that they believe the policy that is optimal given $\omega$ wins, then voters always vote informatively, implying that $\mathbf{p}^{*}=(\omega,\omega)$ can indeed be equilibrium. However, there always coexists another equilibrium, in which candidates again ignore their information. If voters are convinced that $p=0$ is chosen, then any voter with $s<0$ will choose $\tilde{\mu}=\infty$, while a voter with $s>0$ chooses $\tilde{\mu}=0$. Therefore, as in the case of $\Delta>1$, a majority always votes for policy 0, and candidates choose $\mathbf{p}^{*}=(0,0)$.
The next proposition is the paper’s main result and formalizes the above intuitions:
@@
If $\Delta>1$, then there exists a unique pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ and candidates choose $\mathbf{p}^{*}=(0,0)$. If instead $\Delta\leq 1$, then there are multiple equilibria:
  1. 1.
    If $\Delta\leq 1$ and voters have trust in the government, then there exists a pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=\pi(s,\mathbf{p},\tilde{\mu}^{*})$ and candidates choose $\mathbf{p}^{*}=(\omega,\omega)$.
  2. 2.
    If $\Delta\leq 1$ and voters have no trust in the government, then there exists a pure strategy equilibrium in which $\kappa^{*}(s,\mathbf{p},\tilde{\mu}^{*})=0$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
@@ 4 Extensions
In this section I informally discuss further extensions that seem relevant. For the sake of brevity and clarity, throughout this section I assume that $\Delta>1$. In Section 4.1 I discuss how conclusions change if candidates are not only office motivated, but care also about welfare. In Section 4.2 I discuss the implications of differences in candidate valence. In Section 4.3, I discuss the implication of misinformation campaigns. Finally, in Section 4.4 I study the implication of costly deviations from Bayesian rationality.
4.1 Policy Motivated Politicians
Suppose candidates not only care about being elected but also value welfare, as captured by (1). Knowing which state has materialized, candidates are aware of which policy maximizes welfare. Thus, one might hypothesize that adding a policy motivation would improve the situation. However, this is only partially correct. On the one hand, if policy motivation is sufficiently important compared to simply being elected to office, there is indeed an equilibrium in which candidates always choose $\mathbf{p}^{*}=(\omega,\omega)$. To see this, assume the spoils of office have value $V$ and that each candidate receives $V+u(p,\omega)$ if elected and $u(p,\omega)$ otherwise.
$\frac{V}{2}-\Delta\geq V-\Delta-\beta\Leftrightarrow\beta\geq\frac{V}{2}.$
If, however, $\beta<\frac{V}{2}$, there remains a unique pure strategy equilibrium with $\mathbf{p}^{*}=(0,0)$.
Thus, if policy motivation is sufficiently strong, the previously identified problems may not occur. However, there is always another equilibrium in pure strategies where candidates continue to ignore their information despite policy motivation. Assume candidate 1 expects candidate 2 to choose $p_{2}=0$. In this case, the unique best response is to also choose $p_{1}=0$, because policy 0 now wins with certainty. Although candidate 1 cannot influence the policy, she can increase her chances of winning. Therefore, despite policy motivations, and even in the absence of any office motivation, there always exists an equilibrium with $\mathbf{p}^{*}=(0,0)$.
4.2 Valence Differences
Next, consider valence differences as in Ansolabehere and Snyder (2000), Groseclose (2001), Aragones and Palfrey (2002), and more recently Denter (2021). Specifically, assume that voters evaluate candidates not only by the policies they offer but by $v_{i}+u(p_{i},\omega)$, where $v_{i}\in\mathbb{R}$ is the valence utility the voter gets from candidate $i$ in the office and $u(p,\omega)$ is policy utility and defined as before in (1). Let $v_{1}>v_{2}$, implying that candidate 1 has a valence advantage. How does introducing valence alter the model’s conclusions? On the one hand, it has no effect on $\tilde{\mu}^{*}(s)$, which remains as described in Lemma 1, because valence is independent of the state $\omega$. However, valence does affect a voter’s choice at the ballot box for a given belief $\pi(s,\mathbf{p},\tilde{\mu})$. If both candidates choose the same policy, candidate 1 always wins because $v_{1}>v_{2}$. Equilibrium in the extended game resembles equilibrium in Ansolabehere and Snyder (2000). If the valence advantage is not too large, candidate 1 will always choose $p_{1}=0$, while candidate 2 may choose any $p_{2}\in\{0,1\}$ without having a chance to win. If the valence advantage is large enough, candidate 1 may choose any $p_{1}\in\{0,1\}$ and still win with certainty. Thus, it is possible that a candidate with a large valence advantage might use this advantage to propose an optimal policy. In fact, if we combine a large valence advantage with policy motivation as described in Section 4.1, then it is a dominant strategy for candidate 1 to choose $p_{1}=\omega$, and this policy indeed wins the election.
4.3 Disinformation Campaigns
The analysis so far has revealed a robust mechanism through which the electoral incentives of candidates are sufficient to prevent the implementation of adequate policies against climate change. However, climate change is also one of the topics where organized misinformation is most prevalent (see, for example, Lewandowsky, 2021). Perhaps the most famous example of such campaigns is that of ExxonMobil, which, despite having very precise knowledge about the causes and consequences of climate change (see Supran, Rahmstorf, and Oreskes, 2023), chose to downplay the phenomenon in their public communications and editorial-style advertisements in The New York Times (see Supran and Oreskes, 2017).
I follow Papanastasiou (2020), Sisak and Denter (2024) and Denter and Ginzburg (2024) to incorporate misinformation into the model. In particular, I assume that each voter receives a fake signal with a probability of $\alpha$, while with the complementary probability of $1-\alpha$ the signal comes from a proper source and is informative. An informative signal is drawn from the distribution $F_{\omega}$ with corresponding density $f_{\omega}$. I assume that $f_{1}(s)/f_{0}(s)$ is an strictly increasing function of $s$, which is the standard MLRP assumption. Note that in the model described in Section 2 with normally distributed signals this is automatically satisfied. Moreover, without loss of generality let $f_{1}(0)=f_{0}(0)$, implying that $s=0$ is an uninformative signal as before. Unlike proper signals, a fake signal is not informative. In particular, I assume that the distribution of fake signals is $G(s)$ which is independent of $\omega$. For simplicity, further assume that $F_{\theta}$ and $G$ have full support on $\mathbb{R}$.
Because the literature has shown that in the presence of fake news motivated reasoning is often regarding the reliability of an information source, voters cannot change the precision of a signal anymore, but have motivated beliefs about the probability that a given signal is fake or proper. In particular, they can choose $\tilde{\alpha}\in[0,1]$ instead of $\tilde{\mu}\in[0,\infty]$.
For a given signal $s$ and a given choice $\tilde{\alpha}$, the belief about the state $\omega$ is
$\pi(s,\mathbf{p},\tilde{\alpha})=\frac{q\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right]}{q\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{1}(s)\right]+(1-q)\left[\tilde{\alpha}g(s)+(1-\tilde{\alpha})f_{0}(s)\right]}.$
Choosing $\hat{\alpha}=0$ means ignoring the chance that the signal could be fake, and hence the belief equals the perfect Bayesian belief when no fake signals exist. To the contrary, setting $\hat{\alpha}=1$ means ignoring the signal completely because it originates from a fake source.
Lemma 2.
The optimal motivated belief about $\alpha$ is
$\hat{\alpha}^{*}(s)=\left\{\begin{array}[]{cl}1&\text{if }s>0\\ \alpha&\text{if }s=0\\ 0&\text{if }s<0\end{array}\right.$
Hence, any bad news about climate change will be interpreted as fake, while any good news will be interpreted as proper and hence credible. This is in line with recent empirical evidence (Thaler, 2024).
What are the consequences for political decision making? The political equilibrium is the same as in Section 3. The reason is that any signal indicating climate change is severe will be ignored, and hence a majority of the electorate will prefer policy 00 in each state.
Proposition 4.
Suppose $\Delta>1$. There exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\alpha}^{*}(s))=0$, where $\tilde{\alpha}^{*}(s)$ follows from Lemma 2, and in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
@@ 4.4 Costly Distortions
In this section, I extend the previous analysis by accounting for the costs that a voter may incur when distorting the issue’s complexity. The goal is twofold: first, introducing costs allows us to analyze how beliefs may be distorted, which is valuable in itself. Second, it enables us to assess the robustness of the earlier results. I begin by considering a fixed cost of distortion, following the approach of other recent papers, such as Levy (2014) and Le Yaouanq (2023). I then explore a continuous cost function, where the magnitude of the distortion affects the costs, in Section 4.4.2. I will concentrate on conditions under which equilibria with inactive candidates continue to exist.
4.4.1 Fixed Costs
First consider that distorting the issue’s complexity implies a fixed cost of $\gamma>0$. That is, each voter now maximizes
$W(s,\mathbf{p},\tilde{\mu})=AU(s,\mathbf{p},\tilde{\mu})-\gamma\cdot\mathbbm{1}_{\tilde{\mu}\neq\mu}.$
As before, $W(s,\mathbf{p},\tilde{\mu})$ strictly decreases in $s$, implying the greatest incentive to distort $\mu$ when $s>0$ has a voter receiving $s\rightarrow\infty$, and when $s<0$ a voter receiving $s\approx 0$. Of course, if $\gamma$ is too large, no voter has an incentive to choose $\tilde{\mu}\neq\mu$.
Lemma 3.
If $\gamma<\gamma^{+}\equiv(1-q)(\beta(1-\kappa)+\Delta-\kappa)$, then there exists $s^{+}>0$ such that
$\tilde{\mu}^{*}=\left\{\begin{array}[]{cl}0&\text{if }s\geq s^{+}\\ \mu&\text{if }s\in[0,s^{+})\end{array}\right.$
However, if $\gamma\geq\gamma^{+}$, then $\tilde{\mu}^{*}=\mu$ for all $s\geq 0$. Moreover, if $\gamma<\gamma^{-}\equiv q(\beta+\Delta-(\beta+1)\kappa)$, then there exists $s^{-}<0$ such that
$\tilde{\mu}^{*}=\left\{\begin{array}[]{cl}\infty&\text{if }s\in[s^{-},0)\\ \mu&\text{if }s<s^{-}\end{array}\right.$
However, if $\gamma\geq\gamma^{-}$, then $\tilde{\mu}^{*}=\mu$ for all $s<0$.
Having established voters’ optimal deviations from Bayesian rationality, we next seek to determine under which conditions inactivity against climate change, despite better information, remains an equilibrium. First, note that if $\gamma\geq\max\left\{\gamma^{-},\gamma^{+}\right\}$, then no voter distorts their beliefs, and all voters interpret information as perfect Bayesians. It follows from Assumption 2 that in this case, information aggregates correctly in an election with $p_{1}\neq p_{2}$, and therefore the conditions for efficient policy platforms in equilibrium are satisfied. If $\gamma<\min\left\{\gamma^{-},\gamma^{+}\right\}$, there will be both voters who received good news ($s<0$) and those who received bad news ($s>0$) who choose $\tilde{\mu}^{*}\neq\mu$, while others will interpret signals correctly.
Proposition 5.
Suppose $\Delta>1$. There exists $\hat{\gamma}>0$ such that iff $\gamma<\hat{\gamma}$, then there exists a unique pure strategy equilibrium in which voters hold pessimistic beliefs about the enacted policy, $\kappa^{*}(s,\mathbf{p}^{*},\tilde{\mu}^{*})=0$, where $\tilde{\mu}^{*}$ follows from Lemma 3, and in which candidates ignore their information about $\omega$ and choose $\mathbf{p}^{*}=(0,0)$.
The proposition shows that our earlier analysis is robust to the introduction of a fixed cost that needs to be paid if a voter chooses $\tilde{\mu}\neq\mu$. The result is of course not very surprising given Proposition 3. However, one thing to note is that the incentives to distort $\mu$ of voters with $s<0$ do not matter for the result. In fact, those voters vote more often informatively with motivated beliefs than without, which is beneficial for choosing the optimal policy.
4.4.2 Quadratic Costs
Next consider the case of a smooth cost of distortion function. We demand a number of properties our cost function should satisfy. It should be convex increasing when we move away from $\mu$ and zero when there is no distortion. Moreover, it should not be symmetric in the sense that $\tilde{\mu}=\mu+\epsilon$ and $\tilde{\mu}=\mu-\epsilon$ should not lead to the same cost. To see why assume $\mu=1$. A symmetric cost function would imply that $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=0}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=2}$. But note that $\tilde{\mu}=0$ implies that the issue is perceived as infinitely incomplex, and any signal $s\neq 0$ will lead to updating to either 0 or 1. At the same time, $\tilde{\mu}=2$ represents only a gradual increase in complexity. If both have the same cost, then downward distortions are generally “cheaper” than upward distortions. We therefore demand that the cost of equal relative distortions upwards and downwards are the same. This implies the cost function is compressed for downwards distortions. More formally, $\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\lambda\cdot\mu}=\left.C(\tilde{\mu},\mu)\right|_{\tilde{\mu}=\mu/\lambda}$ for all $\lambda>0$ and all $\mu>0$. In particular, we assume the following cost function:
$C(\tilde{\mu},\mu)=\left\{\begin{array}[]{ll}\frac{c}{2}\left(\tilde{\mu}-\mu\right)^{2}&\text{if }\tilde{\mu}\geq\mu\\ \frac{c}{2}\left(\frac{\mu}{\tilde{\mu}}\left(\tilde{\mu}-\mu\right)\right)^{2}&\text{if }\tilde{\mu}<\mu\end{array}\right.$ (6)

### v5

@@ 3.3 Equilibrium Policy Platforms
We can now determine equilibrium platform choices as functions of the parameters of the game. If severe climate change leads to catastrophic baseline welfare level losses, $\Delta\geq 1$, then voters ignore any $s>0$. Our analysis reveals that then a majority always votes for policy 0, whenever $p_{1}\neq p_{2}$. Proposition 2 tells us that in this case candidates never campaign on policy $p=1$, independent of the information they hold. Hence, if severe climate change is catastrophic, then candidates will ignore any information and always choose the optimal policy for mild climate change, $\mathbf{p}^{*}=(0,0)$.
When $\Delta<1$, the set of equilibria becomes larger. The reason is that now voters’ beliefs about the enacted policy become self-fulfilling. If voters trust the government in the sense that they believe the policy that is optimal given $\omega$ wins, then voters always vote informatively, implying that $\mathbf{p}^{*}=(\omega,\omega)$ is indeed an equilibrium. However, there always coexists the equilibrium in which which candidates ignore their information and choose $\mathbf{p}^{*}=(0,0)$.
The next proposition states the paper’s main result and formalizes the above intuitions:
@@
If $\Delta\geq 1$, then there exists a unique pure strategy equilibrium in which $\kappa^{*}=0$ for all $s$ and candidates choose $\mathbf{p}^{*}=(0,0)$. If instead $\Delta<1$, then there exist multiple equilibria. In particular, there exists one pure strategy equilibrium in which, for each $s$, $\kappa^{*}=\pi(s,\mathbf{p},\infty)$ and candidates choose $\mathbf{p}^{*}=(\omega,\omega)$, and there exists another pure strategy equilibrium in which $\kappa^{*}=0$ for all $s$ and candidates choose $\mathbf{p}^{*}=(0,0)$.
The analysis shows that when motivated beliefs are present, a self-fulfilling efficient equilibrium exists as long as the worst consequences of climate change are not too severe, i.e., $\Delta<1$. In this equilibrium, each voter takes her signal at face value and therefore votes informatively. As a result, candidates have the correct incentives and campaign on the policy that is optimal given $\omega$. Consequently, the probability that the winning policy matches $\omega$ is 1.
However, there also always exists a bury-your-head-in-the-sand equilibrium, independent of $\Delta$. In this equilibrium, voters choose to ignore any signal $s>0$, while treating any signal $s<0$ as conclusive evidence that $\omega=0$. The ex ante probability that the correct policy is chosen in this equilibrium is 1q1-q.
4 Discussion and Implications
In this section, we discuss two important implications of the model: that the efficiency of policy-making in the face of climate change depends on both rhetoric and trust in government.
4.1 Trust in Government
When anticipated welfare losses from severe climate change are sufficiently large, the model admits a unique, inefficient equilibrium. By contrast, when these losses are moderate, multiple equilibria arise. In this case, equilibrium selection depends critically on voters’ expectations about how policy choices relate to the underlying state of the world, which we interpret as trust in government.
Following Cabral, (2005) and Besley and Dray, (2024), we define trust as voters’ expectations that political institutions implement policies that are responsive to the true state. In the model, voters hold two distinct beliefs: a belief $\pi$ about the state of the world and a belief $\kappa$ about the probability that policy 1 will be implemented. In the efficient equilibrium with moderate welfare losses, voters’ beliefs about the state take extreme values. As shown in Section 3, motivated belief formation implies that $\pi\in\{0,1\}$, with $\pi=1$ when the voter’s signal satisfies $s>0$ and $\pi=0$ when $s<0$.
In this environment, trusting the political process means that voters expect policy choices to mirror the realized state. Since policy 1 is optimal if and only if climate change is severe, a voter who believes $\pi=1$ expects policy 1 to be implemented with probability one, whereas a voter who believes $\pi=0$ expects policy 1 to be implemented with probability zero. It therefore follows directly in this equilibrium that the perceived probability that policy 1 is implemented coincides with the belief that the state is severe, so that $\kappa=\pi$.
When welfare losses from severe climate change are moderate and voters have trust, they optimally take their signals at face value and vote informatively. Anticipating this behavior, politicians have incentives to follow their information and propose welfare-maximizing policies, thereby sustaining the efficient equilibrium. Trust thus operates as a coordination device: by aligning voters’ expectations about policy implementation with their beliefs about the state, it ensures that information is processed and aggregated efficiently.
Conversely, when trust is low, voters expect policy choices to be weakly related or unrelated to the underlying state, so that $\kappa$ becomes insensitive to information contained in signals. Such pessimistic expectations weaken incentives to process information accurately and can become self-confirming. Anticipating that voters will not respond informatively, politicians optimally ignore their own information and campaign on policies designed for mild climate change, even when climate change is severe. In this way, low trust sustains an inefficient equilibrium.
The model therefore implies a positive relationship between trust in government and the likelihood that electoral competition leads to effective climate policy, provided that welfare losses from severe climate change are not too large. This implication is consistent with Figure 1 and recent empirical evidence showing that higher trust in political institutions is associated with greater provision of public goods, including climate policy (Bose,, 2023).
Importantly, this relationship arises only through the interaction between trust and motivated belief formation; absent motivated reasoning, voters’ beliefs about the state would be independent of expectations about policy implementation, and trust would not affect equilibrium policy choices.
4.2 The Importance of Rhetoric
A first implication concerns political rhetoric. In the model, an efficient equilibrium is less likely to exist when the anticipated welfare losses from severe climate change are sufficiently large. The reason is that larger anticipated losses increase anticipatory disutility and thereby strengthen incentives for voters to engage in defensive belief formation, which undermines coordination on the efficient outcome.
An indirect corollary is that the way climate risks are described may matter for equilibrium selection. To the extent that rhetoric affects voters’ expectations about the magnitude of potential welfare losses from climate change, it can shift the strength of anticipatory disutility and, through this channel, the extent of belief distortion. In this sense, highly alarming descriptions need not translate into greater support for policy action and may instead trigger defensive responses, echoing the concern raised by Clayton et al., (2017).
This argument does not require taking a stand on any particular model of communication. One interpretation is that rhetoric operates through framing, in the sense that equivalent information can induce systematically different perceptions depending on how outcomes are described (Tversky and Kahneman,, 1981, 1989; Ellingsen et al.,, 2012; Spiegler,, 2014). Another interpretation is that rhetoric is informative, in line with cheap-talk and persuasion frameworks in which language conveys signals about states or consequences (Crawford and Sobel,, 1982; Kartik and Van Weelden,, 2019; Kamenica and Gentzkow,, 2011; Alonso and Câmara,, 2016). The present model highlights a simple channel common to both interpretations: by shaping perceived severity, rhetoric can affect anticipatory utility and thus the mapping from information to policy outcomes.
@@ 5 Extensions
In this section we informally discuss further extensions that seem relevant. For the sake of brevity and clarity, throughout this section we assume that $\Delta>1$. In Section 5.1 we study the implication of costly deviations from Bayesian information processing. In Section 5.2 we discuss how conclusions change if candidates are not only office motivated, but care also about welfare.
@@ 5.1 Costly Distortions
In this section, we extend the previous analysis by accounting for the costs that a voter may incur when distorting the issue’s complexity. Introducing costs allows us to scrutinize the above predictions about how beliefs may be distorted, and it enables us to assess the robustness of the earlier results on policy choices.
We distinguish between two types of costs. First, we analyze utility costs from making bad decisions (Bénabou and Tirole,, 2002, 2016; Brunnermeier and Parker,, 2005). We then study purely cognitive costs from choosing distorted beliefs (Levy, (2014); Le Yaouanq, (2023); Engelmann et al., (2024)). For proofs of the statements, see Appendix B.
5.1.1 Cost from Imperfect Decision Making
Assume each voter places a weight $\lambda\in[0,1]$ on anticipatory utility and $1-\lambda$ on the expected utility loss relative to the Bayesian-optimal action. In particular, voters aim at maximizing $W^{UC}=\lambda AU+(1-\lambda)C^{U}$, where $C^{U}\leq 0$ is the utility cost of deviating from Bayesian information processing.
[…]
If the distortion does change behavior, it has an expected cost. This cost is the difference between expected utility under Bayesian beliefs with the optimal action and the expected utility under Bayesian belief with the suboptimal action. The only voters for whom distortion may change the action are those with sufficiently high signals, $s>\tilde{s}$.
When deciding if and how to distort beliefs, a voter faces two potentially optimal choices: (i) $\tilde{\mu}=0$, and thus interpreting any signal as irrelevant; and (ii) choosing the smallest possible $\tilde{\mu}$ such that the choice at the ballot is not altered but anticipatory utility increases. [^5] This option strictly dominates $\tilde{\mu}=\mu$. The voter’s welfare from (i) is
$W^{UC}(\tilde{\mu}=0)=-q(b+\Delta)\lambda-\left(b-\frac{(1-q)(1+b)}{1-\big(1-e^{\frac{2s\mu}{\sigma^{2}}}\big)q}\right)(1-\lambda).$
To determine the utility from (ii), note that the belief about $\omega$ given this must equal $1/(1+b)$. Thus,
$W^{UC}(\tilde{\mu}=\mu_{c})=-\frac{(b+\Delta)\lambda}{1+b}.$
When comparing these two welfare levels, note that when $s$ is such that a Bayesian voter is indifferent between voting for policy 0 or policy $1$, the utility cost from distortion is zero, and thus $W^{UC}(\tilde{\mu}=0)>W^{UC}(\tilde{\mu}=\mu_{c})$. Moreover, $W^{UC}(\tilde{\mu}=0)$ is strictly decreasing in $s$. The reason is that if a voter has a very strong signal, the Bayesian belief is very strong, and the utility cost from acting against this belief is high. Therefore, if $W^{UC}(\tilde{\mu}=0)>W^{UC}(\tilde{\mu}=0)$ even when $s\to\infty$, then any voter with $s>0$ will choose $\tilde{\mu}=0$. This is true if $\lambda$ is sufficiently large. In particular, if
$\lambda>\frac{b(1+b)}{b\big(2-q+(1-q)b\big)+\big(1-q-qb\big)\Delta}\in(0,1).$
Otherwise, for lower levels of $\lambda$, a voter with a very strong signal will choose the non-consequential distortion $\mu_{c}$ and vote according to the signal. Any other voter with $s>0$, however, chooses $\tilde{\mu}=0$, thereby ignoring the signal and voting for policy 0. Hence, the extended model of motivated beliefs provides a more nuanced prediction of voter behavior. If $\lambda$ is large, introducing costs changes nothing. However, if $\lambda$ is sufficiently small, then a strong signal will be interpreted as evidence of severe climate change, and voters receiving such a signal will vote for policy 1. Denote the threshold for a sufficiently strong signal by $\hat{s}$ and note that $\hat{s}>\tilde{s}$.
To shift the equilibrium toward one in which politicians offer policies for severe climate change, this is not yet sufficient. The reason is that if $\mu$ is small—i.e., if climate change is a highly complex issue—then only few voters receive $s>\tilde{s}$. Thus, in this case, there exists a unique equilibrium in which no politician ever offers policy 1. However, for large enough $\mu$, more than 50% of voters receive $s>\tilde{s}$ in state 1, and thus a majority votes for policy 1. This creates incentives for politicians to follow their information and offer policies targeting severe climate change.
5.1.2 Cognitive Costs
Next consider pure cognitive costs from distorting information processing. Distorted information processing can lead to cognitive costs because, for example, sustaining self-serving or overly optimistic beliefs may conflict with one’s deeper desire for accuracy and coherence, creating dissonance and reducing overall well-being. Hence, in the following we assume that there is a cost of $C^{C}(\tilde{\mu},\mu)$ associated with choosing $\tilde{\mu}$.
[…]
$C^{C}(\tilde{\mu},\mu)=\left\{\begin{array}[]{ll}\frac{c}{2}\left(\tilde{\mu}-\mu\right)^{2}&\text{if }\tilde{\mu}\geq\mu\\ \frac{c}{2}\left(\frac{\mu}{\tilde{\mu}}\left(\tilde{\mu}-\mu\right)\right)^{2}&\text{if }\tilde{\mu}<\mu\end{array}\right$. (6)